OpenStudy (anonymous):

Calculus Help!! If f′(x)=5xe^(x^2) and f(0)=0, then f(x)=

OpenStudy (dumbcow):

Use substitution u = x^2 du = 2x dx

OpenStudy (anonymous):

punch what you have into the shape of U'.e^U which integrates into e^U with f'(x) = (5/2).2x.e^(x^2) giving f(x) = (5/2).e^(x^2) + c sub initial value to get at c.

OpenStudy (anonymous):

that's what i did too....i put 2.5 and it is not correct!!

OpenStudy (dumbcow):

what did you put for the constant --> 2.5e^0 +c = 0

OpenStudy (anonymous):

-2.5 is not correct either

OpenStudy (anonymous):

should i put 0?is the answer 0?

OpenStudy (dumbcow):

no -2.5 should be correct ??

OpenStudy (anonymous):

it is incorrect for -2.5!!

OpenStudy (anonymous):

but it shows incorrect for it.stupid online assignment!!thanks anyways:)